OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of two x plus three, over five.?

OpenStudy (anonymous):

just to double check, you mean:  ?

OpenStudy (anonymous):

yes and f−1(x) = The quantity of five x plus three, over two f−1(x) = The quantity of five x minus three, over two f−1(x) = Five over the quantity of two x plus three f−1(x) = Five over the quantity of three minus two x are the options @LifeEngineer

OpenStudy (anonymous):

f−1(x) = The quantity of 5x plus 4, divided by 3. f−1(x) = The quantity of 5x minus 4, divided by 3. f−1(x) = The quantity of negative 3x minus 4, divided by 5. f−1(x) = The quantity of 4 minus 3x, divided by 5.

OpenStudy (anonymous):

well the easiest way to find an inverse of a function is to swap x and y, then solve for y in this case you have y = (2x+3)/5 now swap x and y to get x = (2y+3)/5 finally just solve for y (eg get it alone by performing operations to both sides of the equation) give it a shot and I can give you more help if you'd like!